OpenStudy (anonymous):

what 1 eighth + 5 eighth =

OpenStudy (anonymous):

6 eighths

OpenStudy (anonymous):

\[\frac{1}{8}+\frac{5}8\]

OpenStudy (anonymous):

6 eighths or 3 fourths

OpenStudy (anonymous):

yes nphuongsun93 that way

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }+\frac{ 5 }{ 8 }=\frac{ 6 }{ 8 }=\frac{ 3 }{ 4}\]